### v1

# Workshop on Mechanistic Interpretability

Abstract
Mechanistic interpretability aims to understand how internal components of modern machine learning models, such as weights, activations, and layers, give rise to the model’s overall behavior. One particularly opaque mechanism is attention: despite its central role in transformer models, its mathematical underpinnings and relationship to concepts like feature polysemanticity, superposition, and model performance remain poorly understood. This paper establishes a novel connection between attention mechanisms and multinomial regression. Specifically, we show that in a fixed multinomial regression setting, optimizing over latent features yields optimal solutions that align with the dynamics induced by attention blocks. In other words, the evolution of representations through a transformer can be interpreted as a trajectory that recovers the optimal features for classification.
1 Introduction
Many recent efforts have aimed to provide understanding of modern transformer and large model architectures. Among these, mechanistic interpretations (Ferrando et al., 2024; Sharkey et al., 2025) such as circuit models (Olah et al., 2020; Ameisen et al., 2025), sparse autoencoders (Cunningham et al., 2023), and transcoders (Dunefsky et al., 2024) have been employed to attribute model behaviors to specific neuronal or architectural properties. Many of these attribution methods draw on graphical models (Lindsey et al., 2025) to link discrete structures across neurons, thereby enabling the construction of sparse or discrete representations within computational hyper-graphs (Owhadi, 2022; Bourdais et al., 2025).
The role of sparse polysemanticity in neuron features (Arora et al., 2018; Olah et al., 2020) and feature superposition (Elhage et al., 2022) supports the paradigm that appropriate sparse encodings yield quantized representations in feature space. However, from a theoretical standpoint, it remains unclear why certain modern architectures, such as transformers, offer significant improvements over multilayer perceptrons (MLPs).
We hypothesize that the remarkable performance of transformer architectures, particularly their ability to leverage sparsity and superposition, stems from their implicit execution of feature discovery in conjunction with multinomial regression. This implicit multi-class identification problem naturally encourages sparsity, since one-hot encodings of categorical (or multicategorical) data are exceedingly sparse.
This paper shows that the transformer architecture arises naturally from an optimization problem aimed at recovering the best features for a given multinomial regression task. This connection between transformer blocks and multinomial regression offers a form of intrinsic interpretability for the transformer architecture (Sharkey et al., 2025), while also providing additional support for the sparse superposition theory of mechanistic interpretability.
2 Theoretical Results
Multinomial logistic regression performs classification using a generalized linear model that minimizes the cross-entropy between its predictions and the given class labels. Alternatively, this procedure can be viewed as simple linear regression on the corresponding logits of latent features. The present setting considers a set of input features (possibly in some latent space) $Z=Z(X)\in\mathbb{R}^{S\times F^{i}}$, along with categorical labels $C\in[0,1]^{S\times F^{o}}$ which are one-hot encoded in the sense that $1^{\intercal}C=\sum_{s=1}^{S}C_{s,f^{o}}=1$ for each feature $f^{o}\in\{1,\dots,F^{o}\}$. Training a model $N(Z,\theta)$ with parameters $\theta\in\mathbb{R}^{F^{o}\times F^{i}}$ to fit the labels $C$ involves minimization of the categorical cross-entropy,
$\min_{Z,\theta}-\mathbb{E}\left\langle C,\log\sigma_{i}(N(Z,\theta))\right\rangle,$ (1)
where $\mathbb{E}=\mathbb{E}_{X\sim\rho}$ denotes expectation on $X$ drawn from some distribution $\rho$, and $\sigma_{i}(A)_{ij}=\left(\sum_{k}\exp{A_{kj}}\right)^{-1}\exp{A_{ij}}$ denotes a softmax applied independently to each sequence [^1] Note that exp denotes the element-wise exponential and $\sum_{i}\sigma_{i}(A)_{ij}=1_{j}$ for each column $j$.. Using the definition of softmax, this expression reduces to the expectation of a difference of convex functions in $N(Z,\theta)$,
$\min_{Z,\theta}L(Z,\theta):=\min_{Z,\theta}\mathbb{E}\left[\left\langle 1,\text{LSE}_{i}(N(Z,\theta))\right\rangle-\left\langle C,N(Z,\theta)\right\rangle\right],$ (2)
where $\text{LSE}_{i}$ is the log-sum-exponential function applied across the sequence axis, independently for each feature (as in sequence-to-sequence applications of attention), and we have used that $1^{\intercal}C=1$.
Consider a two stage optimization algorithm that first minimizes over the features $Z$ at each point $X$, followed by optimization over the parameters $\theta$. The goal of this work is to determine the optimal features $Z$ for a fixed $\theta$. To this end, denote the point-wise cross-entropy as
$\ell(Z,\theta)=\left\langle 1,\text{LSE}_{i}(N(Z,\theta))\right\rangle-\left\langle C,N(Z,\theta)\right\rangle.$ (3)
Determining features $Z$ by minimizing this quantity at each point $X$ greedily reduces the global cross-entropy (2).
Consider a linear model $N(Z,\theta)=Z\theta^{\intercal}$. Recalling that the LSE is a smooth approximation of the maximum function, its gradient is a softmax, e.g., (Blanchard et al., 2021), which smoothly approximates the argmax function. Following Appendix A.1, the gradient of the point-wise cross-entropy with respect to the features $Z$ takes the form
$\begin{split}\partial_{Z}\ell(Z,\theta)&=1^{\intercal}\partial_{Z}\text{LSE}_{i}(Z\theta^{\intercal})-C\theta\\ &=\sigma_{i}(Z\theta^{\intercal})\theta-C\theta\\ &=\texttt{CA}\left(Z,\theta\right)-C\theta,\end{split}$ (4)
where $\texttt{CA}(Z,\theta)=\sigma(Z\theta^{\intercal})\theta=1^{\intercal}\partial_{Z}\text{LSE}(Z\theta^{\intercal})$ is the usual cross-attention between $Z$ and $\theta$. In particular, a necessary condition on features $Z=Z(\theta)$ which minimize the point-wise cross-entropy is given by equating this expression to zero, yielding
$\texttt{CA}(Z,\theta)=\sigma_{i}(Z\theta^{\intercal})\theta=C\theta.$ (5)
This occurs exactly when the cross-attention between $Z$ and $\theta$ equals the $\theta$-weighted labels, suggesting that the optimal output of a cross-attention layer is provided when attention uses $\theta$ as key-value pairs.
Let $Z_{0}(X)$ represent an initial guess for the features at each point XX. [^2] This mapping can be injective to ensure a unique representation for $X$. Consider the gradient flow of $\ell(Z,\theta)$ with respect to $Z$, which generates trajectories satisfying the matrix differential equation
$\begin{split}\dot{Z}(t)&=-\partial_{Z}\ell(Z(t),\theta)=C\theta-\texttt{CA}(Z(t),\theta),\\ Z(0)&=Z_{0}(X).\end{split}$ (6)
In particular, it is guaranteed that solutions to this equation will dissipate (point-wise) cross-entropy in continuous time, since
$\dot{\ell}(Z,\theta)=\left\langle\partial_{Z}\ell(Z,\theta),\dot{Z}\right\rangle=-\frac{1}{2}\left\|\partial_{Z}\ell(Z,\theta)\right\|^{2}\leq 0.$ (7)
It follows that running the gradient flow simultaneously for all $X$ will minimize the global cross-entropy $L$ over time. Applying a Strang splitting (Strang, 1968; McLachlan & Quispel, 2002) to this system of ordinary differential equations now yields a system of discrete equations starting from initial condition $Z^{(0)}=Z_{0}(X)$,
$\begin{split}Z^{(\ell+\frac{1}{2})}&=Z^{(\ell)}-\texttt{CA}\left(Z,\theta^{(\ell)}\right)\\ Z^{(\ell+1)}&=Z^{(\ell+\frac{1}{2})}+C\theta^{(\ell+\frac{1}{2})}.\end{split}$ (8)
Remarkably, this mirrors the transformer block in (Vaswani, 2017) using cross-attention. Particularly, the first line of Equation (8) performs cross-attention with a residual connection, while the second line performs a linear mapping of cross-attention’s keys/values (i.e., multiplication by $C$), followed by another residual connection. Note that the keys/values are not a function of the queries $Z$ in this case. Treating Equations 8 as layer-wise propagation of feature information, the parameters $\theta^{(\ell)}$ and $\theta^{(\ell+\frac{1}{2})}$ can then be optimized via gradient descent (or any other optimizer of choice).
Since the use of a linear model $N(Z,\theta)=Z\theta^{\intercal}$ for multinomial logistic regression has connections to cross-attention, it is interesting to consider a quadratic model $N(Z,\theta)=Z\theta Z^{\intercal}$ where $\theta\in\mathbb{R}^{F^{i}\times F^{i}}$ is square. For simplicity, suppose that $\theta=\theta^{\intercal}$ is symmetric and positive definite, so that $\theta=\phi\phi^{\intercal}$ for some lower-triangular square root $\phi\in\mathbb{R}^{F^{i}\times F^{i}}$. Note that in this case, the class target $C\in\mathbb{R}^{S\times S}$ is now a square matrix. Using Equation 3 and repeating the steps above (see Corollary 2 in Appendix A.1 for a derivation), the gradient of the point-wise cross entropy with respect to features $Z$ becomes
$\begin{split}\partial_{Z}\ell(Z,\theta)&=2\sigma_{i}(Z\theta Z^{\intercal})Z\theta-(C+C^{\intercal})Z\theta\\ &=2\,\texttt{SA}(Z\phi)\phi^{\intercal}-(C+C^{\intercal})Z\theta,\end{split}$ (9)
where $\texttt{SA}(X)=\sigma_{i}(XX^{\intercal})X$ is the usual self-attention. Applying the same gradient flow of point-wise cross-entropy, followed by a Strang splitting, we arrive at the corresponding discrete equations, starting from initial conditions $Z^{(0)}=Z_{0}(X)$,
$\begin{split}Z^{(\ell+\frac{1}{2})}&=Z^{(\ell)}-2\,\texttt{SA}\left(Z\phi^{(\ell)}\right){\phi^{(\ell)}}^{\intercal}\\ Z^{(\ell+1)}&=Z^{(\ell+\frac{1}{2})}+(C+C^{\intercal})Z^{(\ell+\frac{1}{2})}\theta^{(\ell+\frac{1}{2})}.\end{split}$ (10)
Again, this structure closely mirrors that of a transformer block with self-attention: the first step performs self-attention followed by a residual connection, while the second step applies a linear transformation.
Returning to the cross-entropy, note that choosing features which minimize the point-wise cross-entropy for a given $\theta$ is sufficient to ensure the minimization over $Z$ in Equation 2. Minimizing over $\theta$ then becomes the training problem typically solved using a variant of stochastic gradient descent.
The implication of these results is that the transformer architecture itself implicitly propagates features through gradient flow on the cross-entropy, connecting nicely to standard multinomial regression. Moreover, it indirectly supports previous observations regarding the roles of sparse coding (Cunningham et al., 2023) and superposition (Elhage et al., 2022) in mechanistic interpretation, since the sparsest features which represent categorical data are one-hot encodings of the class labels. Note that multi-head attention should not substantially change this analysis, directly allowing for multi-class labels in superposition; this will be verified in future work.
3 Conclusions and Limitations
The present formulation applies to deep networks, in contrast to the shallow networks (2 or 3 layers) commonly used in mechanistic interpretability. In particular, this formulation offers a potential explanation for why sparse feature representation can span multiple attention heads and layers, addressing a question raised in (Sharkey et al., 2025).
The primary limitation of this analysis lies in the restriction of model form, where the same set of optimal weights $\theta$ are used across all blocks of the transformer. Additionally, this analysis is restricted to a minimal model which omits the the MLP layer typically present in transformer blocks. We also do not account for the effects of normalization layers, which play a significant role in practical transformer architectures. Notably, normalization shares similarities with whitening, as used in independent component analysis in place of multinomial regression. Alternatively, layer normalization can be interpreted as constraining features to a spherical domain, effectively imposing a directional constraint on the classification problem. This perspective aligns with the “features as directions” view that underlies feature superpositions, as discussed in (Elhage et al., 2022).
Moreover, this analysis assumes that the true class labels $C$ are available during training. An interesting direction for future work is to relax this assumption by introducing an oracle (possibly learned) that provides the labels, potentially integrating with the attention mechanisms within a self-supervised framework. If the oracle is parameterized (or emulated) via an MLP, this begins to bridge the gap between our theoretical formulation and modern state-of-the-art architectures.
Finally, while this analysis offers an interpretation of the underlying mechanics of attention and transformer architectures, modern machine learning models are highly complex and often incorporate additional components beyond the scope of this analysis. Likewise, other state-of-the-art architectures, such as diffusion models, graph neural networks, and state-space models, fall outside the boundaries of this work and present opportunities for future investigation. Ultimately, we hope that the ideas presented here will prove fruitful for understanding attention through the lens of gradient-based dynamics and inspire additional work in this area.
Acknowledgments and Disclosure of Funding
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. SAND#2025-10835O. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
Appendix A Technical Appendices and Supplementary Material
Here we record some computations necessary for fully appreciating the results in the main body.
A.1 Gradients Involving Cross-Entropy and the Log-Sum-Exp Function
The gradient of the cross-entropy $L$ requires differentiating the vector-valued function $\text{LSE}_{i}(Z\theta^{\intercal})$. This can be described succinctly in terms of the operation $A\operatorname{\bar{\otimes}}B$ on matrices $A,B$, which denotes the order-3 tensor with components $(A\operatorname{\bar{\otimes}}B)_{ijk}=A_{ij}B_{ik}$. Note that the row dimensions of $A,B$ must match in order for this diagonal extraction to make sense.
Theorem 1.
The $Z$-derivative of log-sum-exp applied to the linear model $Z\theta^{\intercal}$ satisfies $\partial_{Z}\mathrm{LSE}_{i}(Z\theta^{\intercal})=\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta$.
Proof.
Since $\text{LSE}_{i}(X\theta^{\intercal})\in\mathbb{R}^{s}$ is vector valued, so is its exterior derivative in $Z$ (currying $\theta$), which computes to
$d_{Z}\text{LSE}_{i}(Z\theta^{\intercal})$ $=d_{Z}\log\left(1^{\intercal}\exp(Z\theta^{\intercal})\right)=1^{\intercal}\frac{\exp(Z\theta^{\intercal})\odot dZ\,\theta^{\intercal}}{1^{\intercal}\exp(Z\theta^{\intercal})}=1^{\intercal}\left(\sigma_{i}(Z\theta^{\intercal})\odot dZ\,\theta^{\intercal}\right)$
$=\left(\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right):dZ,$
where $\odot$ denotes element-wise multiplication and the last equality has used that
$\sum_{i,k}\sigma_{i}(Z\theta^{\intercal})_{ij}dZ_{ik}\theta_{jk}=\sum_{i,k}\left[\sigma_{i}(Z\theta^{\intercal})_{ij}\theta_{jk}\right]dZ_{ik}=\sum_{j,k}\left[\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right]_{ijk}dZ_{jk}.$
Since the exterior derivative $df(Z)\in\mathbb{R}^{s}$ of a function $f:\mathbb{R}^{s\times f}\to\mathbb{R}^{s}$ has the expression $\nabla f(Z):dZ$ in terms of the gradient $\nabla f(Z)\in\mathbb{R}^{s\times s\times f}$ at $Z$, the conclusion follows. ∎
Differentiating the term $\mathbb{E}\left\langle 1,\text{LSE}_{i}(X\theta^{\intercal})\right\rangle$ appearing in the cross-entropy $L$ is now as simple as observing that
$\partial_{Z}\left\langle 1,\text{LSE}_{i}(Z\theta^{\intercal})\right\rangle=1^{\intercal}\left(\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right)=\sigma(Z\theta^{\intercal})\theta=\texttt{CA}(Z,\theta).$
Alternatively, there is the following direct calculation for the gradient of the cross-entropy which short-circuits the more elaborate computation in Theorem 1. This will be computed for a general differentiable model $N(Z,\theta)$ and relies on the following simple Lemma.
Lemma 1.
The gradient of the scalar function $f(X)=\mathrm{tr}\left(\mathrm{LSE}_{i}(X)\right)=\langle 1,\mathrm{LSE}_{i}(X)\rangle$ is given by $\nabla f(X)=\sigma_{i}(X)$.
Proof.
Noting that the exterior derivative of $\text{LSE}_{i}$ is
$d\,\text{LSE}_{i}(X)=d\log\left(1^{\intercal}\exp(X)\right)=\frac{1^{\intercal}(\exp(X)\odot dX)}{1^{\intercal}\exp(X)}=1^{\intercal}\left(\sigma_{i}(X)\odot dX\right),$
the exterior derivative of $f$ computes to
$df(X)$ $=\left\langle 1,d\,\text{LSE}_{i}(X)\right\rangle=\left\langle 1,1^{\intercal}\left(\sigma_{i}(X)\odot dX\right)\right\rangle=\left\langle 11^{\intercal},\sigma_{i}(X)\odot dX\right\rangle$
$=\left\langle\sigma_{i}(X),dX\right\rangle=\left\langle\nabla f(X),dX\right\rangle,$
since $\langle A,B\odot C\rangle=\langle A\odot B,C\rangle$, directly yielding the desired gradient. ∎
This provides a way to compute the $Z$-gradient of the cross-entropy for any desired model $N(Z,\theta)$.
Theorem 2.
The $Z$-gradient of the cross entropy $L(Z,\theta)$ satisfies $\partial_{Z}L(Z,\theta)=\mathbb{E}\left[\left(\sigma_{i}(N(Z,\theta))-C\right):\partial_{Z}N(Z,\theta)\right]$ in terms of the model $Z$-gradient $\partial_{Z}N(Z,\theta)$.
Proof.
Recalling that $L(Z,\theta)=\mathbb{E}\left[\left\langle 1,\text{LSE}_{i}\left(N(Z,\theta)\right)\right\rangle-\left\langle C,N(Z,\theta)\right\rangle\right],$ we work term-by-term. The second term under the expectation has the $Z$-exterior derivative
$d_{Z}\langle C,N(Z,\theta)\rangle=\langle C,\partial_{Z}N(Z,\theta):dZ\rangle=\langle C:\partial_{Z}N(Z,\theta),dZ\rangle,$
implying that $\partial_{Z}\langle C,N(Z,\theta)\rangle=C:\partial_{Z}N(Z,\theta)$. Computing the same for the first term in view of Lemma 1 yields
$d_{Z}\left\langle 1,\text{LSE}_{i}\left(N(Z,\theta)\right)\right\rangle$ $=\left\langle\sigma_{i}(N(X,\theta)),dN(Z,\theta)\right\rangle=\left\langle\sigma_{i}(N(X,\theta)),\partial_{Z}N(Z,\theta):dZ\right\rangle$
$=\left\langle\sigma_{i}(N(X,\theta)):\partial_{Z}N(Z,\theta),dZ\right\rangle,$
establishing the gradient $\partial_{Z}\mathrm{LSE}_{i}\left(N(Z,\theta)\right)=\sigma_{i}(N(X,\theta)):\partial_{Z}N(Z,\theta).$ Combining this with the vanishing of the operator commutator $\partial_{Z},\mathbb{E}]=0$ directly yields the conclusion. ∎
Theorem 2 now provides a recipe for computing derivatives of the cross-entropy under the linear and quadratic models considered in this paper.
Corollary 1.
The $Z$-gradient of the cross entropy $L(Z,\theta)$ for a linear model $N(Z,\theta)=Z\theta^{\intercal}$ satisfies $\partial_{Z}L(Z,\theta)=\mathbb{E}\left[\texttt{CA}\left(Z,\theta\right)-C\theta\right].$
Proof.
Observe that the $Z$-gradient of $N(Z,\theta)=Z\theta^{\intercal}$ is given by $\left[\partial_{Z}(Z\theta^{\intercal})\right]_{ijkl}=\delta_{ik}\theta_{jl}$ since
$d_{Z}N(Z,\theta)_{ij}=(dZ\,\theta^{\intercal})_{ij}=\sum_{k}dZ_{ik}\theta_{jk}=\sum_{k,l}\delta_{il}\theta_{jk}dZ_{lk}.$
It follows that $M:\partial_{Z}(Z\theta^{\intercal})=M\theta$ for any matrix $M$, and therefore Theorem 2 directly implies
$\partial_{Z}L(Z,\theta)=\mathbb{E}\left[\left(\sigma_{i}(Z\theta^{\intercal})-C\right):\partial_{Z}(Z\theta^{\intercal})\right]=\mathbb{E}\left[\sigma_{i}(Z\theta^{\intercal})\theta-C\theta\right],$
where $\texttt{CA}(Z,\theta)=\sigma_{i}(Z\theta^{\intercal})\theta$ as claimed. ∎
Corollary 2.
The $Z$-gradient of the cross entropy $L(Z,\theta)$ for a quadratic model $N(Z,\theta)=Z\theta Z^{\intercal}$ with symmetric $\theta$ satisfies $\partial_{Z}L(Z,\theta)=\mathbb{E}\left[2\sigma_{i}(Z\theta Z^{\intercal})Z\theta-(C+C^{\intercal})Z\theta\right].$
Proof.
Observe that the $Z$-gradient of $N(Z,\theta)=Z\theta Z^{\intercal}$ is given by $\left[\partial_{Z}(Z\theta Z^{\intercal})\right]_{ijkl}=\delta_{ik}(Z\theta^{\intercal})_{jl}+\delta_{jk}(Z\theta)_{il}$ since
$d_{Z}N(Z,\theta)_{ij}$ $=(dZ\,\theta Z^{\intercal})_{ij}+(Z\theta\,dZ^{\intercal})_{ij}=\sum_{k}dZ_{ik}(\theta Z^{\intercal})_{kj}+\sum_{k}(Z\theta)_{ik}dZ_{jk}$
$=\sum_{k,l}\left(\delta_{il}(\theta Z^{\intercal})_{kj}+\delta_{jl}(Z\theta)_{ik}\right)dZ_{lk}=\sum_{k,l}\left(\delta_{ik}(Z\theta^{\intercal})_{jl}+\delta_{jk}(Z\theta)_{il}\right)dZ_{kl}.$
It follows that $M:\partial_{Z}(Z\theta Z^{\intercal})=MZ\theta^{\intercal}+M^{\intercal}Z\theta$ for any matrix $M$, and therefore applying Theorem 2 yields
$\partial_{Z}L(Z,\theta)$ $=\mathbb{E}\left[\left(\sigma_{i}(Z\theta Z^{\intercal})-C\right):\partial_{Z}(Z\theta Z^{\intercal})\right]$
$=\mathbb{E}\left[2\sigma_{i}(Z\theta Z^{\intercal})Z\theta^{\intercal}-CZ\theta^{\intercal}-C^{\intercal}Z\theta\right],$
which becomes the stated gradient when $\theta=\theta^{\intercal}$ since $\sigma_{i}(Z\theta Z^{\intercal})^{\intercal}=\sigma_{j}(Z\theta Z^{\intercal})=\sigma_{i}(Z\theta Z^{\intercal})$. ∎